# Vaskar Chakma † School of Artificial Intelligence and Computer Science

### v1

@@ 4.3 Handling Missing Values and Categorical Variables
In this section, two crucial data preprocessing tasks have been focused on and shown in Fig. 6: handling missing values and encoding categorical variables, both of which are vital steps to ensure that the dataset is suitable for machine learning models. Missing values are a common issue in many real-world datasets. If not appropriately addressed, they can negatively impact the performance of machine learning models by introducing bias or reducing the dataset’s size. We chose median imputation as the strategy to handle missing values in the dataset. Median imputation involves replacing missing values with the median value of a column [41]. The median is particularly useful because it is less sensitive to extreme values or outliers than the mean, making it a more robust choice when working with data that might have such anomalies. For example, in the ECG dataset, some numerical columns, such as the ‘rr_interval‘ or ‘qrs_duration‘, may contain missing values for various reasons, such as data collection issues or measurement errors. Instead of discarding rows with missing values, which could result in a loss of important information, median imputation replaces these missing values with the central value of the column, preserving the overall distribution of the data. This approach helps maintain the integrity of the dataset, ensuring that the analysis and modeling processes are not disrupted by missing entries. We used Scikit-learn’s ‘SimpleImputer‘ with the ’median’ strategy to perform this imputation across all relevant numerical columns in the dataset [42].
@@
Figure 6: It displays the organized dataframe containing 800,035 ECG records and 28 clinically relevant features after preprocessing and highlights preserved temporal measurements, electrical axis values, and metadata for machine learning applications.
@@
Once the missing values were handled, the next step was to address the categorical variables in the dataset. Many machine learning algorithms require numerical inputs, so categorical data, often represented as text labels, needs to be converted into a numerical format. In this case, the dataset contains several categorical columns, such as ‘report_0‘, ‘report_1‘, ‘report_2‘, ‘report_3‘, ‘report_4‘, ‘report_5‘, ‘report_6‘, ‘filtering‘, and ‘wct_label‘, all of which contain textual labels that represent different categories or classifications of the data. Label encoding was chosen to convert these categorical text values into numerical labels. Label encoding assigns a unique integer to each category within a given column. For instance, if the ‘report_0‘ column contains the values ‘Normal’, ‘Abnormal’, and ‘Pending’, the Label Encoder would transform these labels into numerical values like 0, 1, and 2. This transformation makes the data usable by machine learning algorithms, which can only process numerical inputs. Scikit-learn’s ‘LabelEncoder‘ has been used for this task, applying it to each categorical column in the dataset [43]. Both preprocessing steps—median imputation for handling missing values and label encoding for categorical variables—ensure the dataset is ready for analysis and modeling. Properly handling missing data avoids biases, maintains the integrity of the dataset, and prevents the loss of valuable information.
@@ 5.1.1 CardioForest: Model Formulation
CardioForest is built upon the Random Forest (RF) framework, enhanced with hyperparameters tuned specifically for ECG feature characteristics and arrhythmic prediction. Random Forest aggregates multiple decision trees trained on bootstrap samples to minimize variance and prevent overfitting [48]. Given a dataset $D$, each tree $t$ is trained on a subset $D_{t}$ sampled with replacement:
@@
where $f_{t}(x)$ is the prediction of the $t$-th decision tree. Each split in a tree aims to minimize the Gini impurity:
@@
where $p_{k}$ is the proportion of samples belonging to a class $k$ at node $N$. Here, $K=2$ for binary classification (WCT vs. non-WCT). CardioForest also incorporates cost-complexity pruning with parameters $\alpha$ to penalize overly complex trees, improving generalization:
@@
where $R(T)$ is the empirical risk (e.g., misclassification rate) of tree $T$.
@@
5.1.2 Comparison Models: GBM, XGBoost, LightGBM
Gradient Boosting Machine (GBM)
GBM constructs an additive model:
$F_{M}(x)=\sum_{m=1}^{M}\gamma_{m}h_{m}(x)$ (5)
$h_{m}(x)\approx-\nabla_{F_{m-1}}L(y,F_{m-1}(x))$ (6)
The model is updated iteratively using a learning rate $\vartheta$:
$F_{m}(x)=F_{m-1}(x)+\vartheta\gamma_{m}h_{m}(x)$ (7)
Extreme Gradient Boosting (XGBoost)
Light Gradient Boosting Machine (LightGBM)
LightGBM accelerates XGBoost’s design using two key strategies:
@@ 5.2 Hyperparameter Tuning for Experimental Setup
To ensure optimal model generalization while preserving clinical relevance, a systematic hyperparameter tuning process [49] was employed across all classifiers (Table 4). Each model underwent a comprehensive grid search procedure, constrained within physiologically plausible and empirically supported parameter ranges.
@@
Table 5: Performance Results Across All Cross-Validation Folds Demonstrating CardioForest’s Superior and Consistent Performance
@@
All hyperparameter tuning outlined in Table 4 was performed using stratified cross-validation, ensuring robust performance estimation under varying data partitions. Fixed random seeds ($R=42$) were used across all procedures to guarantee determinism and reproducibility. Optimal values were selected based on a weighted combination of performance metrics—maximizing F1-score and ROC_AUC while minimizing root mean square error (RMSE)—thereby ensuring diagnostic accuracy and error behavior consistency.
@@ 5.3 Performance Metrics Overview
The performance evaluation (Table 5) of various models through 10 CV [50] revealed that all the classifiers performed well, four machine learning models—CardioForest, XGBoost, LightGBM, and Gradient Boosting—were compared, but CardioForest stood out as the most reliable and consistent for WCT detection. Several metrics were recorded: Accuracy, Balanced Accuracy, Precision, Recall, F1-Score, ROC_AUC, Root Mean Squared Error (RMSE), and Mean Absolute Error (MAE). CardioForest was superior to all the other models in almost all the folds throughout, with an average accuracy of almost 95%, high balanced accuracy ($\sim 0.88$), good precision (greater than 0.93), and good recall ($\sim 0.77$$0.81$). Its ROC_AUC scores were highly significant, indicating excellent classification ability, and its RMSE and MAE scores remained the lowest for all the models, reflecting high overall stability and prediction accuracy. Conversely, XGBoost performed fairly well but with a clear deterioration compared to CardioForest. Average accuracy ranged from 88%–89%, whereas balanced accuracy ranged from 0.71 to 0.73. Precision remained strong (approximately 0.87–0.91), although recall values were significantly lower ($\sim 0.43$$0.48$), demonstrating that the model performed worse at identifying positive cases. Values for RMSE and MAE were higher, indicating higher prediction errors. LightGBM did the worst of all. It had accuracy scores of 83%–85%, with balanced accuracy below 0.66 on average. Precision and recall were lower compared to the rest of the models, which resulted in lower F1-scores and lower ROC_AUC scores. RMSE and MAE were also highest across all models, indicating that LightGBM’s predictive ability on this data was weaker. Gradient Boosting performed both well and poorly. It possessed some of the highest accuracy levels (up to 95.6%) in some folds but also showed instability, particularly in folds 3, 5, and 9, where its performance became extremely poor. Its precision and recall values jumped widely between folds, affecting global stability. Still, Gradient Boosting maintained high ROC_AUC scores, showing a perfect trade-off between sensitivity and specificity when performance was consistent.

### v2

@@ 4.4 Handling Missing Values and Categorical Variables
In this section, two crucial data preprocessing tasks have been focused on and shown in Fig. 6: handling missing values and encoding categorical variables, both of which are vital steps to ensure that the dataset is suitable for machine learning models. Missing values are a common issue in many real-world datasets. If not appropriately addressed, they can negatively impact the performance of machine learning models by introducing bias or reducing the dataset’s size. We chose median imputation [68] as the strategy to handle missing values in the dataset. Median imputation involves replacing missing values with the median value of a column [43]. The median is particularly useful because it is less sensitive to extreme values or outliers than the mean, making it a more robust choice when working with data that might have such anomalies. For example, in the ECG dataset, some numerical columns, such as the ‘rr_interval‘ or ‘qrs_duration‘, may contain missing values for various reasons, such as data collection issues or measurement errors. Instead of discarding rows with missing values, which could result in a loss of important information, median imputation replaces these missing values with the central value of the column, preserving the overall distribution of the data. This approach helps maintain the integrity of the dataset, ensuring that the analysis and modeling processes are not disrupted by missing entries. We used Scikit-learn’s ‘SimpleImputer‘ with the ’median’ strategy to perform this imputation across all relevant numerical columns in the dataset [44].
@@
Figure 6: Organized dataframe structure showing the format of ECG records with 28 clinically relevant features after preprocessing. The dataframe architecture represents the structure used for highlighting preserved temporal measurements, electrical axis values, and metadata for machine learning applications.
@@
Once the missing values were handled, the next step was to address the categorical variables in the dataset. Many machine learning algorithms require numerical inputs, so categorical data, often represented as text labels, needs to be converted into a numerical format. In this case, the dataset contains several categorical columns, such as ‘report_0‘, ‘report_1‘, ‘report_2‘, ‘report_3‘, ‘report_4‘, ‘report_5‘, ‘report_6‘, ‘filtering‘, and ‘wct_label‘, all of which contain textual labels that represent different categories or classifications of the data. Label encoding [69] was chosen to convert these categorical text values into numerical labels. Label encoding assigns a unique integer to each category within a given column. For instance, if the ‘report_0‘ column contains the values ‘Normal’, ‘Abnormal’, and ‘Pending’, the Label Encoder would transform these labels into numerical values like 0, 1, and 2. This transformation makes the data usable by machine learning algorithms, which can only process numerical inputs. Scikit-learn’s ‘LabelEncoder‘ has been used for this task, applying it to each categorical column in the dataset [45]. Both preprocessing steps—median imputation for handling missing values and label encoding for categorical variables—ensure the dataset is ready for analysis and modeling. The complete preprocessing procedure is formalized in Algorithm 1.
Algorithm 1 ECG Data Preprocessing Pipeline
0: Raw MIMIC-IV-ECG dataset $D_{raw}$
0: Preprocessed dataset $D_{processed}$
1: Remove duplicate records using subject_id and study_id
2: Correct biologically implausible values using interpolation
3: Convert timestamps to uniform format
4:for each numerical column $c$ in $D_{raw}$ do
5:  Apply median imputation to handle missing values
6:  $c\leftarrow\text{median}(c)$ for missing entries
7:end for
8:for each categorical column $c$ in $D_{raw}$ do
9:  Apply label encoding: $c\leftarrow\text{LabelEncoder}(c)$
10:end for
11: Perform feature correlation analysis
12: Remove redundant features with correlation >0.95>0.95
13: Apply PCA for dimensionality reduction
14: Standardize features: $X\leftarrow\frac{X-\mu}{\sigma}$
15:return $D_{processed}$
@@ 5.1 Model Selection and Justification
Algorithm 2 CardioForest Training Algorithm
0: Training dataset $D=\{(x_{i},y_{i})\}_{i=1}^{n}$, number of trees $T$, max depth $d_{max}$
0: Trained CardioForest model $\mathcal{F}$
1: Initialize empty ensemble $\mathcal{F}=\{\}$
2: Set hyperparameters: $T=1000$, $d_{max}=20$, $m_{features}=0.6$, $\alpha=0.01$
3:for $t=1$ to $T$ do
4:  Create bootstrap sample: $D_{t}\sim\text{Bootstrap}(D)$
5:  Initialize decision tree $f_{t}$
6:  Build tree using recursive splitting:
7:  while stopping criteria not met do
8:   Select random subset of $m$ features
9:   Find best split minimizing Gini impurity:
10:    $G(N)=1-\sum_{k=1}^{K}p_{k}^{2}$
11:   Split node based on best feature-threshold pair
12:  end while
13:  Apply cost-complexity pruning with $\alpha$:
14:   $R_{\alpha}(f_{t})=R(f_{t})+\alpha|\text{leaves}(f_{t})|$
15:  Add pruned tree to ensemble: $\mathcal{F}\leftarrow\mathcal{F}\cup\{f_{t}\}$
16:end for
17:return $\mathcal{F}$
@@ 5.1.1 CardioForest: Model Formulation
CardioForest is built upon the Random Forest (RF) framework, enhanced with hyperparameters tuned specifically for ECG feature characteristics and arrhythmic prediction [55]. The complete training procedure is described in Algorithm 2. Random Forest aggregates multiple decision trees trained on bootstrap samples to minimize variance and prevent overfitting [51]. Given a dataset $D$, each tree $t$ is trained on a subset $D_{t}$ sampled with replacement:
@@
where $f_{t}(x)$ is the prediction of the $t$-th decision tree. Each split in a tree aims to minimize the Gini impurity [56]:
@@
where $p_{k}$ is the proportion of samples belonging to a class $k$ at node $N$. Here, $K=2$ for binary classification (WCT vs. non-WCT). CardioForest also incorporates cost-complexity pruning [70] with parameters $\alpha$ to penalize overly complex trees, improving generalization:
@@
where $R(T)$ is the empirical risk (e.g., misclassification rate) of tree $T$. The complete CardioForest training process is detailed in Algorithm 2.
5.1.2 Comparison Models: GBM, XGBoost, LightGBM
Gradient Boosting Machine (GBM) [57]
GBM constructs an additive model:
$F_{M}(x)=\sum_{m=1}^{M}\gamma_{m}h_{m}(x)$ (5)
$h_{m}(x)\approx-\nabla_{F_{m-1}}L(y,F_{m-1}(x))$ (6)
The model is updated iteratively using a learning rate $\vartheta$:
$F_{m}(x)=F_{m-1}(x)+\vartheta\gamma_{m}h_{m}(x)$ (7)
Extreme Gradient Boosting (XGBoost) [58]
Light Gradient Boosting Machine (LightGBM) [59]
LightGBM accelerates XGBoost’s design using two key strategies:
@@ 5.1.2 Comparison Models: GBM, XGBoost, LightGBM
Algorithm 3 Stratified Grid Search for Hyperparameter Tuning
0: Dataset $D$, parameter grid $\Theta$, number of folds $K=10$
0: Optimal parameters $\theta^{*}$
1: Initialize best score: $score_{best}=0$
2: Perform stratified K-fold split on $D$
3:for each parameter combination $\theta$ in $\Theta$ do
4:  Initialize fold scores: $scores=[]$
5:  for fold $k=1$ to $K$ do
6:   Split data: $D_{train}^{(k)}$, $D_{val}^{(k)}$
7:   Train model $\mathcal{M}_{\theta}$ on $D_{train}^{(k)}$
8:   Evaluate on $D_{val}^{(k)}$:
9:    $s_{k}=\text{F1-score}(\mathcal{M}_{\theta},D_{val}^{(k)})+\text{ROC-AUC}(\mathcal{M}_{\theta},D_{val}^{(k)})$
10:   $scores\leftarrow scores\cup\{s_{k}\}$
11:  end for
12:  $score_{avg}=\frac{1}{K}\sum_{k=1}^{K}s_{k}$
13:  if $score_{avg}>score_{best}$ then
14:   $score_{best}\leftarrow score_{avg}$
15:   $\theta^{*}\leftarrow\theta$
16:  end if
17:end for
18:return $\theta^{*}$
@@ 5.2 Hyperparameter Tuning for Experimental Setup
To ensure optimal model generalization while preserving clinical relevance, a systematic hyperparameter tuning process [52] was employed across all classifiers (Table 4). Each model underwent a comprehensive grid search procedure, constrained within physiologically plausible and empirically supported parameter ranges [62], as formalized in Algorithm 3.
@@
Table 5: Performance Results Across All 10 Cross-Validation Folds Demonstrating CardioForest’s Superior and Consistent Performance. CardioForest achieved a mean accuracy of 95.19% ($\pm$0.33%) with exceptional stability (CV = 0.35%), substantially outperforming XGBoost (88.44%), LightGBM (84.33%), and GradientBoosting (92.49%).
@@
All hyperparameter tuning outlined in Table 4 was performed using stratified cross-validation, ensuring robust performance estimation under varying data partitions. Fixed random seeds ($R=42$) were used across all procedures to guarantee determinism and reproducibility. Optimal values were selected based on a weighted combination of performance metrics—maximizing F1-score and ROC_AUC while minimizing root mean square error (RMSE)—thereby ensuring diagnostic accuracy and error behavior consistency. The complete optimization procedure is detailed in Algorithm 3.
@@ 5.3 Performance Metrics Overview
The performance evaluation (Table 5) of various models through 10 CV [53, 63] revealed that all the classifiers performed well, four machine learning models—CardioForest, XGBoost, LightGBM, and Gradient Boosting—were compared, but CardioForest stood out as the most reliable and consistent for WCT detection. Several metrics were recorded: Accuracy, Balanced Accuracy, Precision, Recall, F1-Score, ROC_AUC [64], Root Mean Squared Error (RMSE), and Mean Absolute Error (MAE). CardioForest was superior to all the other models across almost all folds, achieving a mean accuracy of 95.19% ($\pm$0.33%), a high balanced accuracy [65] of 88.76% ($\pm$0.79%), an excellent precision of 95.26% ($\pm$0.56%), and a good recall of 78.42% ($\pm$1.57%). Its ROC-AUC scores were highly significant at 0.8886 ($\pm$0.0096), indicating excellent classification ability, and its RMSE (0.2532) and MAE (0.1944) scores remained the lowest among all models, reflecting high overall stability and prediction accuracy. Conversely, XGBoost performed fairly well but with a clear deterioration compared to CardioForest. Average accuracy ranged from 88%–89%, whereas balanced accuracy ranged from 0.71 to 0.73. Precision remained strong (approximately 0.87–0.91), although recall values were significantly lower ($\sim 0.43$$0.48$), demonstrating that the model performed worse at identifying positive cases. Values for RMSE and MAE were higher, indicating higher prediction errors. LightGBM did the worst of all. It had accuracy scores of 83%–85%, with balanced accuracy below 0.66 on average. Precision and recall were lower compared to the rest of the models, which resulted in lower F1-scores and lower ROC_AUC scores. RMSE and MAE were also highest across all models, indicating that LightGBM’s predictive ability on this data was weaker. Gradient Boosting performed both well and poorly. It possessed some of the highest accuracy levels (up to 95.6%) in some folds but also showed instability, particularly in folds 3, 5, and 9, where its performance became extremely poor. Its precision and recall values jumped widely between folds, affecting global stability. Still, Gradient Boosting maintained high ROC_AUC scores, showing a perfect trade-off between sensitivity and specificity when performance was consistent. Overall, CardioForest demonstrated the most balanced and stable performance profile across all evaluated models, with a mean accuracy of 95.19%, an F1-score of 86.02%, and an ROC-AUC of 0.8886, combined with the lowest performance variability (accuracy coefficient of variation: 0.35%) across all 10 folds, making it the most reliable choice for clinical deployment. The prediction procedure with integrated explainability is formalized in Algorithm 4.